# ExploRLLM: Guiding Exploration in Reinforcement Learning with Large Language Models

### v3

@@ VI-A Simulation Setup
The proposed method is trained and evaluated on a simulated tabletop pick-and-place task, as shown in Figure 2. Similar to [21] and [26], we use a UR5e, and the input observation is a top-down RGB-D image. Inspired by [26], we increased the task difficulty by replacing simple blocks with various objects, such as letters. We assess our method in two tasks: a short-horizon (SH) task, “Pick the [pick_letter] and place it in the [place_color] bowl”, and a long-horizon (LH) task, “Put all letters in the bowl of the corresponding color”, as shown in Figure  6(a). In the SH task, each episode starts with three letters and three bowls randomly placed on the table, with pick-and-place actions generated from random language commands. The task is completed when the robot accurately places the chosen letter in the specified bowl. In the LH task, all letters and bowls are randomly arranged, and the task is completed when each letter is placed in a bowl that matches its color.
@@ VI-B Real-World Setup
We validated our approach on a Franka Panda robot equipped with a suction gripper and an RGB-D camera, as shown in Figure 6(a), implementing our policy and code in the EAGERx [27] framework. Given the potential risks to hardware and the time-intensive nature of direct training, we completed training in simulation, with real-robot applications limited to evaluation. We used ViLD for bounding box identification based on object names. To simulate real-world conditions more accurately, we introduced noise to the bounding box center’s position during the training phase in the simulation, mimicking the positional uncertainty inherent in VLM detection. We also added noise to bounding box positions and image inputs, simulating VLM detection uncertainty and camera noise, including lighting variations.
@@ VII Results
Figure 5: Short-horizon ExploRLLM policies can be used in long-horizon tasks with zero-shot LLM planners, e.g. [4].
@@ VII-A Simulation Results
We investigated the effect of varying LLM-based exploration frequencies on training convergence, using $\epsilon\in\{0.0,0.1,\dots,0.9\}$, as shown in Figure 4. An $\epsilon$ of 0 corresponds to stardard SAC. Training was conducted with six random seeds per frequency, and each session began with a 20,000-step warm-up phase without LLM exploration, as no significant policy improvements were observed during this phase. Post-warm-up results, shown in Figure 4 and detailed in Table II for both short- and long-horizon tasks, indicate that ExploRLLM consistently outperforms LLM-only policies across various exploration frequencies.
@@
In the short-horizon task (Figure 4(a)), training without LLM-based exploration is often unstable, resulting in either a successful policy or failure to converge. When the exploration frequency is within $0<\epsilon\leq 0.5$, training stabilizes and converges more quickly, with minimal variation across different $\epsilon$ values. However, increasing $\epsilon$ beyond 0.5 reduces the proportion of online data, slowing progress and introducing greater instability into the training. For long-horizon tasks, Figure 4(b) shows that higher frequencies of LLM-based exploration ($0<\epsilon\leq 0.5$) correlate with faster training. These results highlight the importance of LLM-based exploration in navigating complex tasks by guiding experience toward the optimal region, mitigating challenges from large observation and action spaces. However, similar to the short-horizon tasks, excessive exploration rates introduce instability and slow convergence.
@@
To evaluate the effectiveness of ExploRLLM, we benchmark its performance against four baselines: ExploRLLM without the LLM-based exploration policy, the CaP-style policy [14] (our exploration policy), Socratic Models [4], and Inner Monologue [12]. Our implementations of Socratic Models and Inner Monologue use ViLD [25] as the object detector and GPT-4 [2] as a multi-step planner. The individual steps are executed by a pre-trained CLIPort [26] model with 500 demonstrations. The key difference between Socratic Models and Inner Monologue is that Inner Monologue features a success detector that can identify mistakes.
@@
During evaluation, the letter colors range from seen to unseen colors. Tasks and initialization methods vary, with “NO” indicating no overlap between the initial positions of letters and bowls, and “AO” allowing overlaps. These configurations assess each method’s robustness in handling complex object relationships.
@@
For short-horizon tasks, as shown in Table I, ExploRLLM maintains stable performance, whereas versions without the exploration policy often fail to converge and exhibit high variance in success rates and low-level errors. Our method surpasses LLM-generated policies in success rates, reduces robot behavior errors, and minimizes the performance gap between NO and AO scenarios, emphasizing the exploration policy’s role in correcting FMs’ inaccuracies. In contrast, CLIPort-based methods struggle with novel scenarios or complex geometric object relationships. For long-horizon tasks, RL agents without LLM-based exploitation fail to converge. As shown in Table I, ExploRLLM outperforms Socratic Models, Inner Monologue, and LLM-generated policies, achieving superior results in long-horizon tasks.
@@ VII-B Real-World Results
We conducted real-world evaluations of ExploRLLM in two scenarios: one replicating all letters from the simulation experiments and another introducing the previously unseen letter ‘C’, with each scenario tested over 15 episodes. The short-horizon ExploRLLM achieved success rates of 66.6% for seen letters and 53.3% for the unseen letter scenario. In comparison, the long-horizon ExploRLLM recorded success rates of 40% for seen letters and 33.3% for unseen letters. Despite the Sim2Real gap, our approach shows promising results without additional real-world training. As the VLM extracts the observation space, the RL agent trained in simulation is less distracted by real-world noise. Figure 6(b) illustrates the adaptability of our method in handling diverse object orientations, understanding logical relationships between objects, and executing long-horizon tasks in real-world settings. However, challenges remain with noise in the color and depth perception of objects, which hampers the RL agent’s ability to manipulate objects. Using a photorealistic simulator with extensive domain randomization is expected to improve performance.
@@ VIII Conclusion and Discussion
In this work, we present ExploRLLM, a method that combines RL with FMs. By using actions informed by LLMs and VLMs to guide exploration, we accelerate RL convergence, demonstrating the benefits of integrating the strengths of both RL and FMs. We evaluated our method on tabletop manipulation tasks, showing superior success rates compared to policies based solely on LLMs and VLMs. ExploRLLM also generalizes to unseen colors, letters, and tasks. Ablation experiments with varying levels of LLM-guided exploration further highlighted its significant role in speeding up convergence. Additionally, we validated the method’s ability to transfer learned policies from simulation to real-world scenarios without additional training through real robot experiments. Currently, our framework focuses on tabletop manipulation, but we plan to extend it to a broader range of robotic manipulation tasks. While the system can correct low-level robotic actions, it struggles with mitigating high-level errors that are less frequent in simulations. Future work will focus on addressing these high-level discrepancies.

### v4

@@ VI-A Simulation Setup
We evaluated the proposed method on a simulated tabletop pick-and-place task, as shown in Figure 2. Similar to [22] and [27], we use a UR5e, and the input observation is a top-down RGB-D image. Inspired by [27], we increased the task difficulty by replacing simple blocks with various objects, such as letters. We assess our method in two tasks: a short-horizon (SH) task, “Pick the [pick_letter] and place it in the [place_color] bowl”, and a long-horizon (LH) task, “Put all letters in the bowl of the corresponding color”, as shown in Figure  6(a). In the SH task, each episode starts with three letters and three bowls randomly placed on the table, with pick-and-place actions generated from random language commands. The task is completed when the robot places the chosen letter in the specified bowl. In the LH task, all letters and bowls are randomly arranged, and the task is completed when each letter is placed in a bowl that matches its color.
@@ VI-B Real-World Setup
We validated our approach on a Franka Panda robot equipped with a suction gripper and an RGB-D camera, as shown in Figure 6(a), implementing our policy and code in the EAGERx [28] framework. Given the potential risks to hardware and the time-intensive nature of direct training, we completed training in simulation, with real-robot applications limited to evaluation. We used ViLD to identify bounding boxes based on object names. To simulate real-world conditions more accurately, we introduced noise to the bounding box center’s position during the training phase in the simulation, mimicking the positional uncertainty inherent in VLM detection. We also added noise to bounding box positions and image inputs, simulating VLM detection uncertainty and camera noise, including lighting variations.
@@ VII Results
Figure 5: Short-horizon ExploRLLM policies can be used in long-horizon tasks with zero-shot LLM planners, e.g., [4].
@@ VII-A Simulation Results
We investigated the effect of varying LLM-based exploration frequencies on training convergence, using $\epsilon\in\{0.0,0.1,\dots,0.9\}$, as shown in Figure 4. An $\epsilon$ of 0 corresponds to standard SAC. We trained the agents with six random seeds per frequency, and each session began with a 20,000-step warm-up phase without LLM exploration, as no significant policy improvements were observed during this phase. Post-warm-up results, shown in Figure 4 and detailed in Table II for both short- and long-horizon tasks, indicate that ExploRLLM consistently outperforms LLM-only policies across various exploration frequencies.
@@
In the short-horizon task (Figure 4(a)), training without LLM-based exploration is often unstable, resulting in either a successful policy or failure to converge within the duration of our experiments. Training stabilizes and converges faster when the exploration frequency is within $0<\epsilon\leq 0.5$, with minimal variation across different $\epsilon$ values. However, increasing $\epsilon$ beyond 0.5 reduces the proportion of online data, slowing progress and introducing greater instability into the training. For long-horizon tasks, Figure 4(b) shows that higher frequencies of LLM-based exploration ($0<\epsilon\leq 0.5$) correlate with faster training. These results highlight the importance of LLM-based exploration in navigating complex tasks by guiding experience toward the optimal region, thereby mitigating challenges from large observation and action spaces. However, similar to the short-horizon tasks, excessive exploration rates introduce instability and fail to converge within the duration of the experiments.
@@
To evaluate the effectiveness of ExploRLLM, we benchmark its performance against four baselines: ExploRLLM without the LLM-based exploration policy, the CaP-style policy [14] (our exploration policy), Socratic Models [4], and Inner Monologue [12]. Our Socratic Models and Inner Monologue implementations use ViLD [26] as the object detector and GPT-4 [2] as a multi-step planner. The individual steps are executed by a pre-trained CLIPort [27] model with 500 demonstrations. The key difference between Socratic Models and Inner Monologue is that Inner Monologue features a success detector that can identify mistakes.
@@
During evaluation, the letter colors range from seen to unseen colors. Tasks and initialization methods vary, with “NO” indicating no overlap between the initial positions of letters and bowls and “AO” allowing overlaps. These configurations assess each method’s robustness in handling complex object relationships.
@@
For short-horizon tasks, as shown in Table I, ExploRLLM maintains stable performance. In contrast, versions without the exploration policy have not all converged and exhibit high variance in success rates and low-level errors. Our method surpasses other methods for LLM-generated policies in success rates, reduces robot behavior errors, and minimizes the performance gap between NO and AO scenarios, emphasizing the exploration policy’s role in correcting FMs’ inaccuracies. In contrast, CLIPort-based methods struggle with novel scenarios or complex geometric object relationships. For long-horizon tasks, RL agents without LLM-based exploration fail to converge within the duration of the experiment. As shown in Table I, ExploRLLM outperforms Socratic Models, Inner Monologue, and LLM-generated policies, achieving superior results in long-horizon tasks.
@@ VII-B Real-World Results
We evaluated ExploRLLM in two real-world scenarios: one replicating all letters from the simulation and another introducing the unseen letter ‘C’. Each scenario was tested over 15 episodes. The short-horizon ExploRLLM achieved success rates of 66.6% for seen letters and 53.3% for the unseen letter scenario. In comparison, the long-horizon ExploRLLM recorded success rates of 40% for seen letters and 33.3% for unseen letters. Despite the sim-to-real gap, our approach shows promising results without additional real-world training. As the VLM extracts the observation space, the RL agent trained in simulation is less distracted by real-world noise. Figure 6(b) illustrates the adaptability of our method in handling diverse object orientations, understanding logical relationships between objects, and executing long-horizon tasks in real-world settings. However, challenges remain with noise in the color and depth perception of objects, which hampers the RL agent’s ability to manipulate objects. Using a photorealistic simulator with extensive domain randomization is expected to improve performance.
@@ VIII Conclusion and Discussion
In this work, we presented ExploRLLM, a method that combines RL with FMs. ExploRLLM accelerates RL convergence by using actions informed by LLMs and VLMs to guide exploration, demonstrating the benefits of integrating the strengths of both RL and FMs. We evaluated our method on tabletop manipulation tasks, showing superior success rates compared to policies based solely on LLMs and VLMs. ExploRLLM also generalizes unseen colors, letters, and tasks better. Ablation experiments with varying levels of LLM-guided exploration indicated that extensive tuning of this parameter is unnecessary as values of $0<\epsilon\leq 0.5$ showed convergence improvements. Additionally, we validated the method’s ability to transfer learned policies from simulation to real-world scenarios without additional training through real robot experiments. Currently, our framework focuses on tabletop manipulation, but we plan to extend it to a broader range of robotic manipulation tasks. While the system can correct low-level robotic actions, it struggles with mitigating high-level errors that are less frequent in simulations. Future work will focus on addressing these high-level discrepancies.
IX Acknowledgments
Research reported in this work was partially or completely facilitated by computational resources and support of the Delft AI Cluster (DAIC) [29] at TU Delft (RRID: SCR_025091), but remains the sole responsibility of the authors, not the DAIC team.